OpenStudy (anonymous):

derivative of xsqrt(2x-3). the answer is 3x-3/(sqrt(2x-3)) i just need to see steps of how to get there

OpenStudy (anonymous):

Use the product rule and the chain rule. If it's easier, you can rewrite \(\sqrt{2x-3}\) as \((2x-3)^\frac{1}{2}\).

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

What result are you getting?

OpenStudy (anonymous):

i got to 9(2x-3)^.5+.5x(2x-3)^-.5(2)

OpenStudy (anonymous):

Can you show steps of how you got that result??

OpenStudy (anonymous):

no 9 in front of that..

OpenStudy (anonymous):

Oh, okay, I see.

OpenStudy (anonymous):

So you have \(\sqrt{2x-3}+\dfrac{x}{\sqrt{2x-3}}\), right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

We can rewrite \(\sqrt{2x-3}\) as \(\dfrac{2x-3}{\sqrt{2x-3}}\).

OpenStudy (anonymous):

ohhh